# Blueberry galaxies up to 200 Mpc and their optical/infrared properties

### v1

@@ Abstract
Context. Dwarf highly star-forming galaxies dominated the early Universe and are considered the main driver of its reionization. Direct observations of these very distant galaxies are limited mainly to rest-frame ultraviolet and visible wavelengths, and as a result, some of their properties are out of reach. Therefore, the study of their local analogs, the Green Pea (GP) and Blueberry (BB) galaxies is still of paramount importance.
Aims. This work aims to expand the number of known BBs to smaller distances and the south equatorial sky. In addition to the already known, this new sample of BBs allows for a statistically significant study of their properties probed by visible and infrared (IR) light.
Methods. We utilize HECATE, a catalog of galaxies up to 200 Mpc, which provides optical/IR photometry and characterization of the sources. To select and analyze the sample, we adopt the Pan-STARSS and SDSS optical photometries, and MPA-JHU analysis for sources having spectroscopic observations. Benefited by the multiwavelength coverage, we perform spectral energy distribution fitting to provide homogeneous star-formation rates and stellar masses. The selected BBs are compared with GPs and other dwarf star-forming galaxies (SFGs).
Results. This work has identified 48 BBs out of which 40 are newly discovered. The nearest BB lies at 19 Mpc while 14 BBs are in the south equatorial sky. The BBs tend to be extremely IR red in both WISE and colors, located in different positions in this diagram compared to typical SFGs. Overall, dwarf SFGs with higher specific star-formation rates tend to have redder IR colors.
Conclusions. Although GPs and BBs share many similarities, the latter are the most intensively star-forming sources of the local Universe among dwarf galaxies. Overall BBs are intrinsically bluer in visible light, redder in the IR, less massive, have higher specific star-formation rates and equivalent widths, lower metallicities, and have the most strongly ionized interstellar medium compared to typical SFGs and GPs.
@@
After the Dark Ages, the early Universe transformed from being mostly opaque to being transparent due to the reionization of its intergalactic medium (IGM) during the reionization era (see e.g., Barkana & Loeb 2001, and references therein) which took place at (e.g., Bouwens et al. 2015; Robertson et al. 2015). It is still an open question what caused this pivotal change, and which sources could produce the copious amounts of the required ionizing radiation. Dwarf highly star-forming galaxies (SFGs) are considered mainly responsible because they were numerous and capable of producing large amounts of ultraviolet (UV) and Lyman continuum (LyC) radiation due to hosting a large stellar population (SP) of massive stars (e.g., Shapiro & Giroux 1987; Loeb & Barkana 2001). Active galactic nuclei (AGN) and quasars could have also contributed substantially to the reionization of the IGM (e.g., Volonteri & Gnedin 2009).
@@
Recently, the James Webb Space Telescope (JWST) revolutionized our view of the young Universe by allowing the direct study of the rest-frame UV and visible light of some very distant galaxies at redshifts (e.g., Harikane et al. 2023; Labbé et al. 2023). However, mainly due to their immense distances, the direct study of these early galaxies through other wavelengths (e.g., X-rays, radio) is beyond our current capabilities. Thus, their local analogs, the Green Pea (GP) and Blueberry (BB) galaxies offer the only alternative solution. In fact, recent JWST observations of high-redshift star-forming dwarf galaxies revealed that they show a remarkable resemblance to known relatively nearby GPs (Schaerer et al. 2022; Rhoads et al. 2023).
@@
Green Peas are dwarf starburst galaxies relatively nearby at redshift (Cardamone et al. 2009). They are characterized by low stellar masses (), compact scales (), sub-solar metallicities (log(O/H)+12 8.1) and high star-formation rates (SFR ). Their distinct green color is mainly due to the strong emission lines. Moreover, GPs show significant LyC radiation escape fractions (; Jaskot & Oey 2013; Izotov et al. 2016a, b) which is not common in the low-redshift Universe. The BBs are also highly star-forming dwarfs which are smaller, have even lower metallicities (log(O/H)+12 ), are closer (), and are considered smaller counterparts to GPs (Yang et al. 2017).
@@
Most homogeneous samples of GPs and BBs (Izotov et al. 2011; McKinney et al. 2019; Jaskot et al. 2019; Brunker et al. 2020; Liu et al. 2022; Ding et al. 2023) were based on the Sloan Digital Sky Survey (SDSS). The sample of Yang et al. (2017) was based on SDSS photometry while the rest took advantage of the spectroscopic catalogs available by SDSS. The work of Liu et al. (2022) was based on the spectroscopic coverage of LAMOST (Wang et al. 1996; Su & Cui 2004), and Keel et al. (2022) has also identified some GPs with the help of HST archival data. Overall, the vast majority of known GP and BB samples are limited to the north part of the equatorial sky and relatively large distances.
@@
Besides the properties of these sources that are known through the visible spectrum, only a few works have studied for instance their X-ray (Svoboda et al. 2019) or radio (Sebastian & Bait 2019; Borkar et al. 2024) emission, but they were limited to a small number of sources. Near-infrared (IR) observations were used by Noeske et al. (2003) to study the photometric structure of nearby compact blue dwarf galaxies (BDGs). The WISE survey (Wright et al. 2010) has also been used for the discovery of BDGs. Griffith et al. (2011) have discovered two sources with extremely red colors in mid-IR wavelengths. Chandola et al. (2023) selected 11 blue compact galaxies with bright mid-IR emission and studied their H I emission.
@@
The majority of the previous studies have focused on high-z samples, but local samples are needed to be able to study in more detail their properties and for multi-wavelength follow-up. An extended and relatively more nearby sample of BBs will allow the statistically significant study of the BBs’ IR and visible emission, and could also serve as a base for future observations and further investigation. This work aims to expand the sample of one of the most extreme classes of SFGs, the BBs, using the Heraklion Extragalactic Catalogue (HECATE; Kovlakas et al. 2021)111https://hecate.ia.forth.gr/. Section 2 describes the selection criteria, and Section 3 shows the resulting sample, the main results regarding their visible/IR properties, and comparisons with other samples of BDGs, GPs, and BBs. A discussion based on the results is presented in Section 4, and Section 5 concludes the article. A Planck Collaboration et al. (2016) cosmology is adopted throughout the paper (, ). Unless stated otherwise, uncertainties and confidence intervals (C.Is) refer to 68%, and the size of the histogram’s bins is based on Scott’s rule (Scott 1979).
2 Selection criteria.
@@
In order to extend the sample of BBs beyond the footprint of SDSS, we performed a search that depended on the photometric rather than the spectroscopic properties of the sources. The selection was based on HECATE which is an all-sky catalog of 204,733 galaxies within redshift ( Mpc) with more than 50% completeness in B-band luminosity up to 170 Mpc. It offers distances, sizes, positions, SFRs, stellar masses, and metallicities, as well as morphological and nuclear activity classifications along with the galaxies’ IR and optical photometry. Thus, it is perfectly suited to selecting specific types of galaxies and examining their properties. We performed our search by using the second, enhanced version of HECATEv2.0 (Kyritsis et al. in prep) which provides additional information like optical photometry from Pan-STARRS (Chambers et al. 2016), and more robust estimations of stellar masses and SFRs that were calculated using the methods of Kouroumpatzakis et al. (2023). HECATEv2.0 provides photometry through Pan-STARRS for the majority of the sources covered by this survey, down to a declination of .
@@
We adopted the Pan-STARRS Kron photometry and the PSF photometry which was also used for the selection since most BBs appear to be spatially compact (Yang et al. 2017). We also adopted the SDSS cModelMag and the Petrosian photometries provided by SDSS. To avoid missing BBs, which have some of the bluest colors found in galaxies, all four aforementioned photometric systems were complementary used. This allowed us to cross-test the results since on many occasions one of the different photometries may vary for various reasons (e.g., contamination from other sources, effects like varying photometric/atmospheric conditions between the separate filter observations, etc.).
@@
In order to quantify their differences and define the final photometric scheme for the selection of the BBs, we compared the colors of the optical photometric schemes mentioned above (Figure 1). This comparison involved all the galaxies in HECATE classified as star-forming (no AGN, composite, or LINERS) with a signal-to-noise ratio higher than three (S/N ) in the g, and r bands in all four Pan-STARRS and SDSS photometries. The best agreement is between the cModelMag and Petrosian magnitudes as SDSS provided both. The standard deviation of the difference between their colors is 0.05 dex. The Pan-STARRS Kron magnitudes show good agreement on average () and the standard deviation of their difference is dex. However, galaxies based on Pan-STARRS Kron magnitudes appear to be redder for blue galaxies and bluer for red ones concerning both cModelMag and Petrosian SDSS photometries. As expected, the Pan-STARRS PSF photometry shows the largest scatter compared to the rest of the photometries. Moreover, it shows on average magnitude bluer colors compared to Kron photometry throughout the color range. The results of the linear regression fits between the colors of the different photometries are summarized in Table 1.
@@
The cModelMag photometry corresponds to an aperture defined by a radial profile that is a linear combination of the best fit of an exponential and a de Vaucouleurs profile. Thus, because cModelMag is defined to encompass most of the light of extended sources like galaxies, it is considered the most trustworthy photometric scheme for this work and, therefore, it is used to define the selection criteria. However, to be able to expand the BB sample to parts of the sky not covered by SDSS, we also take into account sources that were measured only by Pan-STARRS.
@@
The selection of the galaxies was based on their photometric properties following the scheme proposed by Yang et al. (2017) who, in order to find BBs, selected sources with , , , and (mag) adopting the SDSS cModelMag photometry. These criteria were defined by simulating the emission of BBs with intensive star formation and strong ionization that leads to large equivalent widths ( Å) and very young SP ages. The simulations modeled galaxies with an old SP component with age equal to 900 Myr, and one very recent starburst episode at 4 Myr with the code Starburst99 (Leitherer et al. 1999).
@@
For this work, we allow slightly more relaxed criteria of mag which corresponds to the scatter of the comparison between the cModelMag and Kron photometries (Table 1), partially covering for photometric uncertainties, or slightly different star formation histories (SFHs) and metallicities from those assumed by Yang et al. (2017). Thus, the selection limit for the SDSS cModelMag and Petrosian photometries is (mag). Also, because Pan-STARRS Kron photometry appears redder for blue galaxies compared to SDSS (Figure 1), we further allow for mag for the selection criteria applied to Pan-STARRS photometry. This value was estimated based on the linear regression fit (Table 1) between SDSS cModelMag and Pan-STARRS Kron photometries applied to the limit (mag). We kept galaxies that fulfill these criteria simultaneously in the SDSS cModelMag and Petrosian photometries, or simultaneously in the Pan-STARRS Kron and PSF photometries. Additionally, we calculated the average colors by combining all the mentioned photometries. Sources with average colors that did not fulfill the criteria applied to Pan-STARRS photometry were removed (3 sources omitted). We also omitted sources with problematic photometry by limiting the accepted colors to and respective uncertainties . The selection criteria are summarized in Equation LABEL:eq:Selection shown in the Appendix A. This filtering process resulted in 83 sources that fulfilled the selection criteria.
@@
Moreover, because we are interested in the IR properties of the galaxies, we omitted sources where the WISE extended source catalog flag star_contamination is true to avoid those with photometries contaminated by bright stars. This was the case for 8 sources (PGC 360, 5817, 6336, 3876, 42031, 45482, 8881, 4094118) all of which are spectroscopically confirmed stars and were removed. The final step involves visual inspection of the selected galaxies. The galaxies that were omitted through visual inspection are: PGC 46329, 48755, 4081747, 46495, 4084118, and 6723779 which are also spectroscopically confirmed stars; PGC 5060407 due to its presence in a crowded field of background and foreground sources; PGC 28757, 6741333, 6741249, 4107113, 6742056, 5059213, 6740710, 40683, and 5057032 which resemble ultra-faint galaxy candidates; PGC 5078149, 774317, 1076594, and 101538 which are blue star-forming regions that are part of significantly larger or spiral galaxies; and PGC 33553, 53905, 49767, 56870, 41771, 5082584, and 46982 which are sources that have strongly contradicting photometry between SDSS and Pan-STARRS.
@@ 3 Results
3.1 Proposed sample of Blueberry galaxies.
This process yielded 48 BBs, the most distant at 198 Mpc and the nearest at 19 Mpc. Table 2 reports the selected galaxies and their , and colors from the different photometries. The sources are reported in ascending order of their weighted color combining all photometries with weights equal to 1.2, 1.1, 1.0, and 0.9 for the SDSS cModelMag, SDSS Petrosian, Pan-STARRS Kron, and Pan-STARRS PSF respectively. Table 3 reports some fundamental characteristics of the galaxies like their , specific SFR (), sizes and distances, along with metallicities and EWs of the , , H, and H emission lines for sources that have spectra and were analyzed by MPA-JHU (Kauffmann et al. 2003; Brinchmann et al. 2004; Tremonti et al. 2004). Tables 2 and 3 are also provided as an online Table that includes all columns’ uncertainties. Figure 2 shows the optical images composed of g, r, and i-band observations of the selected sources as imaged by SDSS, or Pan-STARRS for sources not covered by SDSS.
The nearest BB found is BB 42/PGC 1005611 at 19 Mpc. The second nearest source is BB 3/PGC 71496 at 75 Mpc, and the third is BB 44/PGC 4080012 at 86 Mpc. Eight sources are common with the Yang et al. (2017) BBs. Those are the BBs PGC 3359860, 3995162, 101527, 4571707, 46812, 139572, and 4120682, 3328781 which are also part of the McKinney et al. (2019) and Jaskot et al. (2019) samples. In the following analysis, for comparisons that include Yang et al. (2017) BBs, we only consider new sources identified by this work.
Although Yang et al. (2017) reports 25 BBs at distances below 200 Mpc () 17 of them are not included in HECATE. However, 15 out of these 25 BBs have distances higher than 175 Mpc where HECATE’s completeness drops significantly. The other 8 galaxies that are included in HECATE have been successfully recognized as BBs by this selection method. The nearest BB reported by Yang et al. (2017) is at 101 Mpc while this work reports 7 galaxies below 100 Mpc. The mode of the distribution of distances of Yang et al. (2017) BBs is Mpc while the mode of the distribution of distances of BBs in this work is Mpc.

### v3

# Blueberry galaxies up to 200 Mpc and their optical and infrared properties

@@ Abstract
Context. Dwarf highly star-forming galaxies (SFGs) dominated the early Universe and are considered the main driver of its reionization. However, direct observations of these distant galaxies are mainly confined to rest-frame ultraviolet and visible light, limiting our understanding of their complete properties. Therefore, it is still paramount to study their local analogs, the green pea (GP) and blueberry (BB) galaxies.
Aims. This work aims to expand our knowledge of BBs by identifying a new sample that is closer and in the southern sky. In addition to the already known BBs, this new sample will allow for a statistically significant study of their properties probed by visible and infrared (IR) light.
Methods. By utilizing the HECATE catalog, which provides optical and IR photometry and characterization of galaxies, along with data from Pan-STARSS and SDSS, this study selects and analyzes a new sample of BBs. We employed spectral energy distribution fitting to derive homogeneous measurements of star-formation rates and stellar masses. Additionally, we measured emission-line fluxes, including , through spectral fitting.
Results. Through this work, we identified 48 BBs, of which 40 were first recognized as such, with the nearest at 19 Mpc. Fourteen of the BBs are in the south sky. The BBs tend to be extremely IR red in both WISE and colors, distinguishing them from typical SFGs. Dwarf SFGs with higher specific star-formation rates tend to have redder IR colors.
Conclusions. Blueberry galaxies stand out as the most intensely star-forming sources in the local Universe among dwarf galaxies. They exhibit unique characteristics, such as being intrinsically bluer in visible light, redder in the infrared, and less massive. They also have higher specific star-formation rates, equivalent widths, lower metallicities, and the most strongly ionized interstellar medium compared to typical SFGs and GPs.
@@
After the Dark Ages, the early Universe transformed from being mostly opaque to being transparent due to the reionization of its intergalactic medium (IGM) during the reionization era (see e.g., Barkana & Loeb 2001, and references therein), which took place at (e.g., Bouwens et al. 2015; Robertson et al. 2015). What caused this pivotal change and which sources could have produced the copious amounts of the required ionizing radiation are still open questions. Dwarf highly star-forming galaxies (SFGs) are mainly considered to be responsible because they were numerous and capable of producing large amounts of ultraviolet (UV) and Lyman continuum (LyC) radiation due to hosting a large stellar population (SP) of massive stars (e.g., Shapiro & Giroux 1987; Loeb & Barkana 2001). Active galactic nuclei (AGNs) and quasars could have also contributed substantially to the reionization of the IGM (e.g., Volonteri & Gnedin 2009).
@@
Recently, the James Webb Space Telescope (JWST) revolutionized our view of the young Universe by allowing the direct study of the rest-frame UV and visible light of some very distant galaxies at redshifts (e.g., Harikane et al. 2023; Labbé et al. 2023). However, mainly due to their immense distances, the direct study of these early galaxies through other wavelengths (e.g., X-rays, radio) is beyond our current capabilities. Thus, their local analogs, namely green pea (GP) and blueberry (BB) galaxies, offer the only alternative solution. In fact, recent JWST observations of high-redshift star-forming dwarf galaxies revealed that they show a remarkable resemblance to known and relatively nearby GPs (Schaerer et al. 2022; Rhoads et al. 2023).
@@
Green peas are dwarf starburst galaxies relatively nearby at a redshift of (Cardamone et al. 2009). They are characterized by low stellar masses (), compact scales (), sub-solar metallicities (log(O/H)+12 8.1), and high star-formation rates (SFRs; i.e., ). Their distinct green color is mainly due to strong emission lines. Moreover, GPs show significant LyC radiation escape fractions (; Jaskot & Oey 2013; Izotov et al. 2016a, b) which is not common in the low-redshift Universe. Blueberry galaxies are also highly star-forming dwarfs, but they are smaller, have even lower metallicities (log(O/H)+12 ), are closer (), and are considered smaller counterparts to GPs (Yang et al. 2017).
@@
Most homogeneous samples of GPs and BBs (Izotov et al. 2011; McKinney et al. 2019; Jaskot et al. 2019; Brunker et al. 2020; Liu et al. 2022; Ding et al. 2023) are based on the Sloan Digital Sky Survey (SDSS). The sample of Yang et al. (2017) was based on SDSS photometry, while the rest took advantage of the spectroscopic catalogs available by SDSS. The work of Liu et al. (2022) was based on the spectroscopic coverage of the Large Sky Area Multi-Object Fibre Spectroscopic Telescope (LAMOST; Wang et al. 1996; Su & Cui 2004), and Keel et al. (2022) has also identified some GPs with the help of Hubble Space Telescope (HST) archival data. Overall, the vast majority of known GP and BB samples are limited to the north part of the equatorial sky and relatively large distances.
@@
Apart from the properties of these sources that are known through the visible spectrum, only a few works have studied, for instance, their X-ray (Svoboda et al. 2019) or radio (Sebastian & Bait 2019; Borkar et al. 2024) emission, but these studies were limited to a small number of sources. Near-infrared observations were used by Noeske et al. (2003) to study the photometric structure of nearby compact blue dwarf galaxies (BDGs). The WISE survey (Wright et al. 2010) has also been used for the discovery of BDGs. Griffith et al. (2011) discovered two sources with extremely red colors in mid-IR wavelengths. Chandola et al. (2023) selected 11 blue compact galaxies with bright mid-IR emission and studied their H I emission, and Lumbreras-Calle et al. (2022) found 466 extreme emission-line galaxies combining the Javalambre Photometric Local Universe Survey (J-PLUS; Cenarro et al. 2019) and the ALLWISE catalog (Cutri et al. 2021).
@@
The majority of the previous studies have focused on high-z samples, but local samples are needed to be able to study the properties of BBs in more detail and for multi-wavelength follow up. An extended and relatively more nearby sample of BBs will allow for a statistically significant study of the BBs’ IR and visible emission and could also serve as a base for future observations and further investigation. This work aims to expand the sample of one of the most extreme classes of SFGs, the BBs, using the Heraklion Extragalactic Catalogue (HECATE; Kovlakas et al. 2021).111https://hecate.ia.forth.gr/ This work is organized as follows: Section 2 describes the selection criteria, and Section 3 shows the resulting sample, the main results regarding their visible and IR properties, and comparisons with other samples of BDGs, GPs, and BBs. A discussion based on the results is presented in Section 4, and Section 5 concludes the article. A Planck Collaboration et al. (2016) cosmology is adopted throughout the paper (, ). Unless stated otherwise, uncertainties and confidence intervals (CIs) refer to 68%, and the size of the histogram’s bins is based on Scott’s rule (Scott 1979).
2 Selection criteria
@@
In order to extend the sample of BBs beyond the footprint of SDSS, we performed a search that depended on the photometric rather than the spectroscopic properties of the sources. The selection was based on HECATE, which is an all-sky catalog of 204,733 galaxies within redshift ( Mpc) with more than 50% completeness in B-band luminosity up to 170 Mpc. It offers distances, sizes, positions, SFRs, stellar masses, and metallicities as well as morphological and nuclear activity classifications along with the galaxies’ IR and optical photometry. Thus, it is perfectly suited to selecting specific types of galaxies and examining their properties. We performed our search using the second enhanced version of HECATE, HECATEv2.0 (Kyritsis et al. in prep), which provides additional information, such as optical photometry from Pan-STARRS (Chambers et al. 2016), and more robust estimations of stellar masses and SFRs that were calculated using the methods of Kouroumpatzakis et al. (2023). HECATEv2.0 provides photometry through Pan-STARRS for the majority of the sources covered by this survey down to a declination of .
@@
We adopted the Pan-STARRS Kron and the point-spread function (PSF) photometry, which was also used for the selection since most BBs appear to be spatially compact (Yang et al. 2017). We also adopted the SDSS cModelMag and the Petrosian photometries provided by SDSS. To avoid missing BBs, which have some of the bluest colors found in galaxies, all four aforementioned photometric systems were used in a complementary manner. This allowed us to cross-test the results since on many occasions one of the different photometries could have varied for different reasons (e.g., contamination from other sources, effects such as varying photometric or atmospheric conditions between the separate filter observations).
@@
In order to quantify their differences and define the final photometric scheme for the selection of the BBs, we compared the colors of the optical photometric schemes mentioned above (Figure 1). This comparison involved all the galaxies classified as star forming in HECATE (no AGNs, composite, or LINERs) with a signal-to-noise ratio higher than three (S/N ) in the g and r bands in all four Pan-STARRS and SDSS photometries. The best agreement was found between the cModelMag and Petrosian magnitudes, as SDSS provides both. The standard deviation of the difference between their colors is 0.05 dex. The Pan-STARRS Kron magnitudes show good agreement on average, , and the standard deviation of their difference is dex. However, galaxies based on Pan-STARRS Kron magnitudes appear to be redder for blue galaxies and bluer for red ones concerning both cModelMag and Petrosian SDSS photometries. As expected, the Pan-STARRS PSF photometry shows the largest scatter compared to the rest of the photometries. Moreover, it shows, on average, magnitude bluer colors compared to Kron photometry throughout the color range. The results of the linear regression fits between the colors of the different photometries are summarized in Table 1.
@@
The cModelMag photometry corresponds to an aperture defined by a radial profile that is a linear combination of the best fit of an exponential and a de Vaucouleurs profile. Thus, because cModelMag is defined to encompass most of the light of extended sources (e.g., galaxies), it is considered the most trustworthy photometric scheme for this work, and therefore, we used it to define the selection criteria. However, to be able to expand the BB sample to parts of the sky not covered by SDSS, we also took into account sources that were measured only by Pan-STARRS.
@@
The selection of the galaxies was based on their photometric properties following the scheme proposed by Yang et al. (2017) who, in order to find BBs, selected sources with , , , and (mag), adopting the SDSS cModelMag photometry. These criteria were defined by simulating the emission of BBs with intensive star formation and strong ionization that lead to large equivalent widths ( Å) and very young SP ages. The simulations modeled galaxies with an old SP component with an age equal to 900 Myr and one very recent starburst episode at 4 Myr with the code Starburst99 (Leitherer et al. 1999).
@@
For this work, we allowed slightly more relaxed criteria of mag, which corresponds to the scatter of the comparison between the cModelMag and Kron photometries (Table 1) and partially compensates for photometric uncertainties or slightly different star formation histories (SFHs) and metallicities from those assumed by Yang et al. (2017). Thus, the selection limit for the SDSS cModelMag and Petrosian photometries is (mag). Also, because Pan-STARRS Kron photometry appears redder for blue galaxies compared to SDSS (Figure 1), we further allowed for mag for the selection criteria applied to Pan-STARRS photometry. This value was estimated based on the linear regression fit (Table 1) between SDSS cModelMag and Pan-STARRS Kron photometries applied to the limit (mag). We kept galaxies that fulfill these criteria simultaneously in the SDSS cModelMag and Petrosian photometries or simultaneously in the Pan-STARRS Kron and PSF photometries. Additionally, we calculated the average colors by combining all the mentioned photometries. Sources with average colors that did not fulfill the criteria applied to Pan-STARRS photometry were removed (three sources were omitted). We also omitted sources with problematic photometry by limiting the accepted colors to and respective uncertainties . The selection criteria are summarized in Equation LABEL:eq:Selection shown in the Appendix A. This filtering process resulted in 83 sources that fulfilled the selection criteria.
@@
Moreover, because we are interested in the IR properties of the galaxies, we omitted sources where the WISE extended source catalog flag star_contamination is true in order to avoid sources with photometries contaminated by bright stars. This was the case for eight sources (PGC 360, 5817, 6336, 3876, 42031, 45482, 8881, and 4094118), all of which are spectroscopically confirmed stars and were removed. The final step involved visual inspection of the selected galaxies. The galaxies that were omitted through visual inspection are PGC 46329, 48755, 4081747, 46495, 4084118, and 6723779, which are also spectroscopically confirmed stars. The reasons for their omission are as follows: PGC 5060407 due to its presence in a crowded field of background and foreground sources; PGC 28757, 6741333, 6741249, 4107113, 6742056, 5059213, 6740710, 40683, and 5057032 because they resemble ultra-faint galaxy candidates; PGC 5078149, 774317, 1076594, and 101538 because they are blue star-forming regions that are part of significantly larger or spiral galaxies; and PGC 33553, 53905, 49767, 56870, 41771, 5082584, and 46982, as they are sources that have strongly contradicting photometry between SDSS and Pan-STARRS.
@@ 3 Results
3.1 Proposed sample of blueberry galaxies
The selection process yielded 48 BBs, with the most distant at 198 Mpc and the nearest at 19 Mpc. Table 2 reports the selected galaxies and their and colors from the different photometries. The sources are reported in ascending order of their weighted color, combining all photometries with weights equal to 1.2, 1.1, 1.0, and 0.9 for the SDSS cModelMag, SDSS Petrosian, Pan-STARRS Kron, and Pan-STARRS PSF, respectively. Table 3 reports some fundamental characteristics of the galaxies, including their , specific SFR (sSFR; ), sizes, and distances along with metallicities and EWs of the , , H, and H emission lines for sources that have spectra and were analyzed by MPA-JHU (Kauffmann et al. 2003; Brinchmann et al. 2004; Tremonti et al. 2004). Uncertainties for Tables 2 and 3 are provided in the online222Tables 2 and 3 are available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (147.231.47.162) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/ version. Figure 2 shows the optical images composed of g, r, and i band observations of the selected sources as imaged by SDSS or Pan-STARRS, for sources not covered by SDSS.
The nearest BB found is BB 42 (PGC 1005611) at 19 Mpc. The second nearest source is BB 3 (PGC 71496) at 75 Mpc, and the third nearest is BB 44 (PGC 4080012) at 86 Mpc. There are eight sources in common with the Yang et al. (2017) BBs. These are the BBs PGC 3359860, 3995162, 101527, 4571707, 46812, 139572, 4120682, and 3328781, and they are also part of the McKinney et al. (2019) and Jaskot et al. (2019) samples. In the following analysis, for comparisons that include Yang et al. (2017) BBs, we only consider new sources identified by this work.
Although Yang et al. (2017) reports 25 BBs at distances below 200 Mpc (), 17 of them are not included in HECATE. However, 15 of these 25 BBs have distances higher than 175 Mpc, where HECATE’s completeness drops significantly. The eight galaxies that are included in HECATE have been successfully recognized as BBs by our selection method. The nearest BB reported by Yang et al. (2017) is at 101 Mpc, while this work reports seven galaxies below 100 Mpc. The mode of the distribution of distances of Yang et al. (2017) BBs is Mpc, while the mode of the distribution of distances of BBs in this work is Mpc.